OpenStudy (anonymous):

an ethernet cable is 4 m long and has a mass of 0.2 kg . a transverse wave is produced by plucking one end of the taut cable. the pulse makes 4 trips down and back along the cable in 0.8 sec , what is the tension in the cable?

OpenStudy (anonymous):

Let "\(T\)" be the tension in the cable, and then the vertical component of the cable tension is: \(F_v = T×Sin(30) \) Now, try to solve it using the formula.. @YamadaTasnim

OpenStudy (anonymous):

\[trips / time= (1/2L) \sqrt{T/(kg/m)} \] i use this formula.. got it right! @Persia

OpenStudy (michele_laino):

the emitted frequency \( \large \nu \) by the cable is given by the subsequent formula: \[\Large \nu = \frac{1}{{2L}}\sqrt {\frac{T}{\rho }} \;\left( {{\text{Hz}}} \right)\] where \( \large \rho \) is the linear density of the cable, and \( \large T \) is the tension which is acting on the cable